Prev Next

SW&IT Meeting Minutes 2020-10-05

Attendees: Carl, David, Igor, Matthieu, Stephen, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20201005
LR20200921 - Bec, Matthieu 80.0
LR20200921 - Holmberg, Carl 80.0
LR20200921 - Edmeades, David 80.0
LR20200921 - Sola, Igor 80.0
LR20200921 - Hooper, Stephen 80.0
LR20200921 - Bryant, Trey 40.0
LR20200921 - Zhang, Yang 80.0
LR20200921 - Zhang, Xianyu 80.0

News from management

  • Paul Grenz half-time 2-3 months

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports

CH

  • Last week:
    • AGW:
      • Met with SB and JPH to plan our way forward on integrating the binary encoder/ACC 84 interface board.
      • Reviewed JPH's post-lab test encoder/interface board software configuration notes.
    • ARGOS: Searched for VirtualBox VM files to back up to Google Drive.
    • DIMM: Reviewed wiki and s/w.
    • DMS: Peter Gray is happy with the iMonnit sensor settings.
    • LBC: 64 bit transition.
      • Attended zoom session to review functionality of LBC tools with OK, YZ, SH. Most LBC issues appear to be addressed.
      • Retested and patched fits verification function, added file settling wait time in RB_Science_py3.py to close out JH’s original Python 3 update request, began a general fits file test script to address limits of my existing file copy test. YZ developed a robust, generalized viewing script.
    • Safety training.

  • This week:
    • AGW: Write an encoder/interface board software integration plan, schedule lab time for testing.
    • ARGOS: Upload VM files to Google Drive.
    • DIMM: System/software walk through with SH.
    • DMS: Record desired sensor settings and process for updating premier license in wiki. Write a small script to email alert a setting change.
    • LBC 64 bit transition
      • Participate in follow up discussions.

DE

  • Mountain synology remote access to support, iterating with them
  • Veeam update bugfix seems to be working
  • Removed VMHost2 from Veeam backup
  • Wolfgang Zoom followup
    • I've set the meetings to require a Zoom account, which he does not have. I feel this is not an unreasonable minimum to require, and it seems like something that Zoom could force us to do as a security enhancement.
  • Cisco training
  • Set up volume @25TB for AO log data in Tucson, but 13TB of existing data is too much to copy over the mountain link. Let's decide on a portable storage device to sneakernet the data down.
  • All backups running smoothly
Next:
  • Juan Carlos Guerra LBT email?
  • Monday YZ VM OS upgrade
  • Plan for Google Docs/Shane sorted? I am under the impression that the task has been set to him but if there is need for a particular file immediately we can take care of that.
  • Storage plan refresh
  • More Cisco
  • RMA HD on its way to me

YZ

  • Previous:
    • Mostly LBC related work:
      • Follow up discussions on LBC's transition to 64-bit machine for operations (including zoom meeting with Olga, Stephen, and Carl). Mostly working, perhaps some small tweaking needed.
      • LBC bias level adjustment with Alessandro and Olga.
      • Spent most of the time looking at the RB_Science program's issues. All appear to be resolved now, will release my own version this week.
    • Some small amount of TMS and AO-Cravat related work.
  • Next:
    • Some remaining LBC work (github reorganization, follow-up discussion, program release etc.)
    • Need to spend most of the time on AO-Cravat.
    • Perhaps dedicate some small amount of time to study and implement the new TMS strategy.

TB

  • Previous:
    • Last changes to report
    • Getting started on Plotly with dummy data
    • Experiment with Plotly chart studio (to help w/ CSS styling)
  • Next:
    • Get farther with Plotly
    • Add last extra figure to report and talk about it (9/26)

SH

  • DIMM testing
  • PMC network (checks)
  • LDAP VPN
  • ROWG stuff
next:
  • LeRoy PEPSI servers
  • LDAP VPN
  • Webapps (add PMS and checkin)
  • DIMM stuff.

MB

  • Last week
    • OC, Management
    • budget 2021-23
    • leadership workshop
    • ROWG meeting
    • ATT mostly
    • ROWG document
    • review different ITs, PMS for VATT
  • This week
    • ATT mostly
    • Argos-calunit project description
    • AGw umac, PBR, AO-MCS, ROWG-ops, TMS, Ao-cravat meetings

IS

Last week:
  • OT
    • ADI (Angular Differential Imaging) (or parallactic rotation) enhancement to OT. It meant adding a new rotation mode to the OT. Position vs Parallactic. This new rotation mode, in principle will only apply to LUCI1/2 and LUCI fraternal scripts. Also expanded the Argos Collimation scripts for a more generic collimation script.
  • PIT
    • Could do some more work on Close for Proposals opening and closing dates, and listing of current proposals per partner to the pit admin tool. Still some more work to do.

This week:
  • Safety Training
  • PIT
    • the Close for Proposals opening and closing dates, and listing of current proposals per partner to the pit admin tool
  • LUCI
    • Try to do the same I did for the TCS VM in a docker container, and finish the to write the doc.
    • Add more debugging information in Journalizer DB storing logic. Trying to figure out why sometimes the mask config file doesn't upload.
Topic revision: r11 - 05 Oct 2020, DavidEdmeades
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback